Suppose you know that a company's stock currently sells for $67 per share and the required return on the stock is 12 percent. You also know that the total return on the stock is evenly divided between a capital gains yield and a dividend yield. If it's the company's policy to always maintain a constant growth rate in its dividends, what is the current dividend per share?
Answer #1

The total retum is evenly divided into the capital gain yield and dividend yield. Hence, the capital gain yield is \(6 \%\) and dividend yield is \(6 \%\).

Calculate the dividend per share as follows:

Expected return = Dividend yield + Capital gain yield

$$ \begin{aligned} &=\frac{D_{0} \times(1+g)}{P_{0}}+g \\ 0.12 &=\frac{D_{0} \times(1+0.06)}{\$ 67}+0.06 \\ D_{0} \times(1+0.06) &=(0.12-0.06) \times \$ 67 \\ D_{0} &=\frac{\$ 4.02}{1.06} \\ &=\$ 3.79 \end{aligned} $$

Therefore, the dividend per share is \(\$ 3.79\)